OpenStudy (anonymous):

4x - y = 8 -8x + 2y = -5

OpenStudy (anonymous):

use system of equations

OpenStudy (anonymous):

4x - y = 8 -8x + 2y = -5 So we need to find a way when we add these equations that something will cancel. So let's multiply the top one by 2. 8x - 2y = 8 -8x + 2y = -5 Added together we get 0x + 0y = 3 Since 0 does not equal 3, we say there is no solution, or give it a null set! :D

OpenStudy (precal):